### v3

@@ Introduction
Through simulation of the beam propagation, we find the most favorable detection plane to be the QCL’s front facet. The system’s point spread function (PSF), measured using a knife-edge technique, shows a better than two-fold improvement in lateral and axial confocal resolution, effectively surpassing the diffraction limit. We attribute this enhanced resolution to the confocal imaging configuration and the characteristics of the QCL which acts as a confocal aperture. The depth of focus of $\lambda/5$, acquired from the phase information, exceeds limitations of confocal imaging systems, resulting in a further significant improvement in axial resolution. We combine the high-resolution capability with a fast beam steering configuration to deliver a 0.5 Mpixel image in less than two minutes. This compares favorably with other reported imagers of both the raster scanning (single pixel) and multipixel focal-plane array based solutions, alike. To accommodate a wide range of applications requiring different scan areas, we also investigate the system’s performance with different focusing objective numerical apertures. Finally, we illustrate the system’s depth of focus resolution by performing 3D tomographic imaging of hidden features of a packaged silicon chip.
@@ Results
Higher NAs further enhance the lateral resolution of LFI imaging systems (Fig. 7a). Inspection of subwavelength features of a resolution target is possible via the incorporation of a $0.5$ NA, presenting noticeable improvement upon the $0.3$ NA used when inspecting the system’s PSF. The modulation transfer function (MTF) is computed to determine the lateral resolution (Fig. 7b), which surpasses the diffraction limit ($71\mu$m or $0.7\lambda$ for $\mathrm{NA}=0.5$, Fig. 7c). This is attributed to the pinhole effect from the QCL’s waveguide upon reinjection, similar to confocal microscopy techniques [39]. Interestingly, the astigmatic behaviour shown in Fig. 5 results in distinct $x$ and $y$ lateral resolutions for $\mathrm{NA}=0.3$, while the highest $0.5$ NA displays identical lateral resolutions in both $x$ and $y$, resulting in overall better imaging performance.
@@ Conclusion
Through beam propagation simulations and experimental characterisation, we measured the system’s point spread function, revealing significant improvements in lateral resolution beyond the diffraction limit and in axial resolution through confocal depth-of-focus control. This enhancement is attributed to the LFI confocal aperture, which achieves a DOF as small as $\lambda/5$ by leveraging phase information.

### v4

@@ Introduction
Through simulation of the beam propagation, we find the most favorable detection plane to be the QCL’s front facet. The system’s point spread function (PSF), measured using a knife-edge technique, shows a better than two-fold improvement in lateral and axial confocal resolution. We attribute this enhanced resolution to the confocal imaging configuration and the characteristics of the QCL which acts as a confocal aperture. The depth of focus of $\lambda/5$, acquired from the phase information, exceeds limitations of confocal imaging systems, resulting in a further significant improvement in axial resolution. We combine the high-resolution capability with a fast beam steering configuration to deliver a 0.5 Mpixel image in less than two minutes. This compares favorably with other reported imagers of both the raster scanning (single pixel) and multipixel focal-plane array based solutions, alike. To accommodate a wide range of applications requiring different scan areas, we also investigate the system’s performance with different focusing objective numerical apertures. Finally, we illustrate the system’s depth of focus resolution by performing 3D tomographic imaging of hidden features of a packaged silicon chip.
@@ Results
Higher NAs further enhance the lateral resolution of LFI imaging systems (Fig. 7a). Inspection of subwavelength features of a resolution target is possible via the incorporation of a $0.5$ NA, presenting noticeable improvement upon the $0.3$ NA used when inspecting the system’s PSF. The modulation transfer function (MTF) is computed to determine the lateral resolution, achieving values of $71\mu$m or $0.7\lambda$ for $\mathrm{NA}=0.5$ (Fig. 7b-c). This is attributed to the pinhole effect from the QCL’s waveguide upon reinjection, similar to confocal microscopy techniques [39]. Interestingly, the astigmatic behaviour shown in Fig. 5 results in distinct $x$ and $y$ lateral resolutions for $\mathrm{NA}=0.3$, while the highest $0.5$ NA displays identical lateral resolutions in both $x$ and $y$, resulting in overall better imaging performance.
@@ Conclusion
Through beam propagation simulations and experimental characterisation, we measured the system’s point spread function, revealing significant improvements in lateral and axial resolution through confocal depth-of-focus control. This enhancement is attributed to the LFI confocal aperture, which achieves a DOF as small as $\lambda/5$ by leveraging phase information.